Find all points on the y-axis that are 6 units from the point (4, -3).
Answer by FrankM(1040)

The y axis has the value 0 for all x, i.e x=0 'is' the y axis. (0,y) describes every point as well.
The distance between two points is similar to the Pythagorean theorem as we take the difference in x and y and solve for the 'third side'.
(x2-x1)^2+(y2-y1)^2=D^2
(4-0)^2+(-3-y)^2=36
16+(-3-y)^2=36
(-3-y)^2=20
9+6y+y^2=20
y^2+6y-11=0
You can solve a number of ways to get the two results.
y=-3+/-2sqr5 or 1.472, -7.472 These are the two points on the y axis.